### v1

# Capturing quantum phase transition in the ultraviolet region by holography

Abstract
We reveal for the first time that ultraviolet (UV) observables can diagnose quantum phase transitions (QPTs). In a class of holographic models exhibiting metal-insulator transitions, we study two types of UV observables – high-frequency conductivity and short-range entanglement. Remarkably, we find that the derivatives of these UV observables exhibit extrema near the quantum critical point. Analytical results show these critical behaviors arise from the deformation of the asymptotic bulk geometry. Moreover, these UV diagnostics show enhanced robustness to thermal fluctuations compared to typical infrared (IR) diagnostics, providing a clean method to identify quantum criticality at finite temperature. This work opens a new window for exploring quantum critical phenomena via UV physics in the laboratory.
Introduction.— Quantum phase transitions (QPTs) are prominent phenomena in strongly correlated systems, conventionally characterized by the critical point separating distinct ground states [1, 2]. Diagnosing the occurrence of the QPT by locating the critical point at the zero temperature is a fundamental problem in quantum critical phenomenon. Usually, the infrared (IR) fixed point of the renormalization group (RG) flow typically dominates their universal characteristics; consequently, traditional studies have primarily focused on IR physics [3, 4]. However, diagnosing QPTs solely through IR observables faces two key challenges. (i) Many IR observables are highly susceptible to thermal smearing, limiting their effectiveness in diagnosing QPTs [5]. (ii) They are insensitive to high-energy features, failing to capture microscopic signatures that may also encode critical information [1].
Recent advances in both theory and experiment provide a novel approach to tackling these challenges. On the theoretical side, certain strongly correlated systems show that high-energy (UV) degrees of freedom influence low-energy (IR) scaling behavior [6, 7], which implies that the information of quantum criticality, typically associated with IR dynamics, may also be encoded in the UV regime. Moreover, the inherently high energy scale of UV physics far exceeds typical thermal fluctuations. This property suggests UV observables are more insensitive to thermal effects, making them powerful candidates for diagnosing QPTs at finite temperature. On the experimental side, techniques such as scanning tunneling spectroscopy and angle-resolved photoemission can detect microscopic UV structures with high thermal stability in strongly correlated systems [8, 9, 10]. Nevertheless, it remains unclear whether the quantum criticality leaves any detectable signatures in the UV region. One major obstacle results from the exponential growth of the computational complexity with the size of the strongly correlated system. At this stage the AdS/CFT correspondence, which is also dubbed as the holographic duality, provides a powerful framework to address this problem [11, 12, 13, 14, 15, 16, 17, 18]. It asserts that a strongly coupled quantum system in D dimensions is equivalent to a weakly coupled gravity theory in $D+1$ dimensional AdS spacetime. This duality maps the complex UV dynamics of the strongly correlated boundary theory to controlled deformations of the asymptotic AdS geometry in the bulk. Therefore, the UV behavior of the system in principle may be revealed by the AdS/CFT correspondence.
In this letter, we show for the first time that UV observables can probe quantum criticality with greater precision and robustness to thermal fluctuations than their IR counterparts. Specifically, we consider a holographic model characterized by metal-insulator phase transitions [2, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. By performing fourth-order UV asymptotic expansions and high-precision numerical simulations, we examine two types of UV observables, including high-frequency conductivity and short-range entanglement measures [29, 30, 31, 32, 33, 34, 35]. We find that the derivatives of high-frequency AC conductivity and short-range entanglement measures exhibit extrema near the quantum critical point (QCP), with critical signatures appearing as subtle variations resolvable at the sixth significant digit. These critical behaviors are analytically shown to arise from the leading deformation of the asymptotic bulk geometry. Moreover, at finite temperatures, UV observables diagnose QPTs more accurately than their IR counterparts, which enables them to serve as reliable probes of quantum criticality. Our findings establish a new paradigm for diagnosing QPTs, highlighting the unexpected richness of the UV region and its potential to complement traditional IR-focused approaches, especially at finite temperature.
Holographic setup and phase structure.— We consider the four-dimensional Einstein-Maxwell-Dilaton-Axion (EMDA) model, a well-established holographic framework for studying metal-insulator transitions (MITs) driven by translational symmetry breaking [20, 36]. The Lagrangian for this model is given by:
$\begin{split}\mathcal{L}=&R+6\,\cosh\,\psi-\frac{3}{2}\left[\left(\partial\psi\right)^{2}+4\,\sinh^{2}\psi\left(\partial\chi\right)^{2}\right]\\ &-\frac{1}{4}\,\cosh^{\gamma/3}\left(3\psi\right)F^{2},\end{split}$ (1)
where $R$ is the Ricci scalar, and $F=dA$ is the Maxwell field strength (with $A$ being the Maxwell field). $\psi$ is a scalar dilaton field, and $\chi$ is an axion field. The presence of the axion, specifically chosen as $\chi=\bar{k}x$, explicitly breaks translation symmetry in the $x$-direction, mimicking the effect of a lattice in condensed matter system. The parameter $\gamma$ governs the coupling between the dilaton and the Maxwell field, playing a crucial role in determining the nature of the IR fixed points.
We seek planar black brane solutions with the following ansatz:
$ds^{2}=\frac{1}{z^{2}}\left(-Pdt^{2}+\frac{dz^{2}}{P}+V_{1}dx^{2}+V_{2}dy^{2}\right),$ (2)
$P=U(1-z)\left(1+z+z^{2}-\frac{\mu^{2}z^{3}}{4}\right),$
$A=\mu(1-z)adt,\qquad\chi=\bar{k}x,\qquad\psi=z^{3-\Delta}\phi,$
where $\mu$ is the chemical potential of the dual field theory, and $\Delta=2$ is the conformal dimension of the scalar operator dual to $\psi$. In this coordinate system, the AdS boundary is located at $z=0$, and the black brane horizon is at $z=1$. The functions $U$, $V_{1}$, $V_{2}$, $a$, and $\phi$ are functions of the radial coordinate $z$ only.
Three dimensionless parameters characterize the solutions: the normalized Hawking temperature $T\equiv\bar{T}/\mu$, the normalized lattice amplitude $\lambda\equiv\bar{\lambda}/\mu$, and the normalized wave number $k\equiv\bar{k}/\mu$. Here, $\bar{\lambda}$ corresponds to the boundary value of the dilaton field $\psi(0)$, effectively setting the strength of the lattice deformation. The chemical potential $\mu$ serves as the scaling unit, placing the system in a grand canonical ensemble. See Appendix A for more details.
The EMDA model exhibits rich phase structure at zero temperature, where varying $\lambda$ or $k$ can drive distinct QPTs by altering the nature of the IR fixed points. Specifically, the IR fixed points can be classified by the value of $\gamma$ under the linear perturbations of the near horizon geometry, which is $AdS_{2}\times\mathbb{R}_{2}$ [20, 36, 37].
To distinguish metallic and insulating phases, we examine the temperature derivative of DC conductivity in the zero temperature limit: $\partial_{T}\sigma_{DC}\big{|}_{T\to 0}<0$ (metallic), $\partial_{T}\sigma_{DC}\big{|}_{T\to 0}>0$ (insulating), and $\partial_{T}\sigma_{DC}\big{|}_{T\to 0}=0$ (QCP). This allows us to construct the phase diagram in the $\lambda$-$k$ plane [36, 37].
For the purpose of this study, we fix the dilaton-Maxwell coupling exponent to $\gamma=-1/6$ and the lattice amplitude to $\lambda=2$. We approach the zero-temperature limit by setting the numerical temperature to $T=10^{-7}$. Under these conditions, we find that the QCP is located at $k_{c_{0}}=0.43$, with the metallic phase for $k>k_{c_{0}}$ and the insulating phase for $k<k_{c_{0}}$. Next we focus on the behavior of UV observables near the QCP.
UV Signatures in High-Frequency Conductivity.— Now we explore the quantum critical signature in the UV region by examining the AC conductivity $\sigma(\omega)$, where the high-frequency limit ($\omega\to\infty$) directly probes the UV behavior of the dual field theory. In this limit, the real part of the AC conductivity approaches $1$ asymptotically, and specifically, we find that $\mathrm{Re}[\sigma_{AC}]$ admits the expansion as $\omega\to\infty$:
$\mathrm{Re}[\sigma_{AC}]=1+\frac{C_{\sigma}}{\omega^{2}}+\frac{P_{\sigma}}{\omega^{4}}+O(\omega^{-6}),$ (3)
where $C_{\sigma}$ remains constant and independent of $k$, while $P_{\sigma}$ provides the leading $k$-dependence at $O(\omega^{-4})$. Since $P_{\sigma}$ appears at the order $\omega^{-4}$, extracting its variation with $k$, which directly reflects the $k$-dependence of $\sigma_{AC}$, requires extremely high numerical precision, and we present all the details on the numerical analysis in Appendix B. Here we remark that in high-frequency region with low temperature, the solutions exhibit rapid oscillations, requiring dense discretization along the radial direction $z$. The numerical convergence is guaranteed through grid refinement tests and residual analysis, with typical discretizations using 1000+ grid points.
Remarkably, our numerical analysis reveals that the third derivative of the UV conductivity with respect to the control parameter $k$, $\partial_{k}^{3}\sigma_{\scriptscriptstyle{UV}}$, exhibits a pronounced extremum near the QCP, as shown in Fig. 1. As the temperature goes down, the location of the extremum goes closer to the QCP. This phenomenon maintains as the frequency becomes higher. This indicates that the critical signature of MIT is encoded as the appearance of the extremum in the high frequency region of the conductivity, indeed.
Figure 1: Third derivatives of $\sigma_{AC}$ with respect to $k$ for various $T$ at $\omega/\mu=20$. The black dashed line represents the critical point.
[…]
From a holographic perspective, this UV sensitivity can be understood through the bulk geometry’s response to the QPT. The near-boundary behavior of the bulk fields determines the high-frequency conductivity, specifically the leading corrections to the asymptotic AdS geometry. As the system approaches the QCP, the entire bulk geometry must reorganize to interpolate between different IR fixed points. This global reorganization necessarily affects the UV region through the equations of motion, which couple different radial scales. The fact that we observe extremal behavior in the third derivative, rather than the conductivity itself, indicates that the UV encodes information about the rate of change of the system’s response to the tuning parameter—a quantity intimately related to the critical susceptibility.
Furthermore, from a renormalization group perspective, while the universal critical exponents are indeed determined by the IR fixed point, the RG flow itself spans all energy scales. The UV region contains information about the initial conditions of this flow. Near a QPT, where the correlation length diverges, even UV observables can become sensitive to which basin of attraction the system belongs to, as the crossover scale between UV and IR behaviors shifts dramatically. This is particularly evident in holographic theories where the bulk geometry provides a geometric realization of the RG flow.
This finding prompts a crucial question: Is this UV sensitivity specific to transport properties, or does it represent a more general feature of quantum criticality? To explore this universality, we next examine whether similar signatures appear in geometric probes of entanglement.
Figure 2: First derivative of HEE with respect to $k$ for various widths at $T=10^{-7}$. The orange diamond markers ($P_{S}$) represent analytical results, while the curves show numerical results. The black dots mark the minimum of each curve, and the black dashed line represents the critical point.
Figure 3: Third derivatives of MI (upper panel) and EWCS (lower panel) with respect to $k$ at $T=10^{-3}$. Dashed curves represent analytical results; solid curves show numerical results. Upper panel: The separation $b=10^{-4}$ is fixed while $a$ and $c$ are varied. Lower panel: $a=c=10^{-2}$ are fixed while $b$ is varied. The black dots indicate the extrema of each curve, and the black dashed line represents the critical point. The small discrepancy between analytical and numerical results arises from finite values of $a,b$, and $c$; this discrepancy diminishes as these parameters decrease, while the extremal behavior remains robust.
UV Quantum Entanglement Signatures of Criticality.— It is well known that quantum information measures, such as the entanglement entropy (EE), mutual information (MI) as well as reflected entropy, are traditionally used to probe long-range correlations in the IR region and thus be able to diagnose QPTs. In holography their function of diagnosing QPT has also been justified in the IR since for large subsystem configurations they exhibit extrema near the quantum critical point (QCP), as shown in Appendix C. In this section we demonstrate that the above quantities surprisingly reveal sharp signatures of the QPT in the UV regime, indeed.
We investigate three key holographic entanglement measures: holographic entanglement entropy (HEE) for a strip of width $w$, MI for two disjoint regions $A$ and $C$ separated by $B$ (with widths $a,b,c$), and entanglement wedge cross-section (EWCS), a holographic measure of reflected entropy.
Strikingly, the first derivative of HEE with respect to the control parameter $k$, $\partial_{k}S$, develops a distinct extremum near the QCP for small widths ($w\sim 10^{-3}$ to $0.3$) (Fig. 2). Similarly, the third derivatives of MI and EWCS, $\partial_{k}^{3}I$ and $\partial_{k}^{3}E_{W}$, also display pronounced extrema near the QCP in their UV limits (Fig. 3). These observations indicate that quantum criticality is robustly imprinted in UV entanglement structures.
Analytically, we can trace the UV behavior of these measures to specific $k$-dependent coefficients in their series expansions. For HEE, we find that it admits the UV expansion as $w\to 0$:
$S(w)=\frac{L_{y}}{4\mu G_{N}}\bigg{[}-\frac{C_{-1}}{w}+P_{S}+O(w)\bigg{]},$ (4)
where $C_{-1}$ remains constant and independent of $k$, while $P_{S}\equiv\hat{V}_{1}^{(1)}$ captures the leading $k$-dependence, where $\hat{V}_{1}^{(1)}=\partial_{z}V_{1}|_{z=0}$ represents the leading deviation of the $g_{xx}$ metric component from pure AdS4 (see Appendix D for details). The first derivative $\partial_{k}S$ isolates this leading $k$-dependent term after $k$-independent contributions cancel. As shown in Fig. 2, $P_{S}(k)$ (orange diamond markers) accurately reproduces the numerical $\partial_{k}S$ for small $w$. For MI and EWCS, their dominant $k$-dependence in the UV arises from a higher-order term, $P_{IE}\equiv\hat{U}^{(3)}-\hat{V}_{1}^{(3)}-2\hat{V}_{2}^{(3)}$, which captures more subtle deviations in the UV geometry. We need the third derivatives $\partial_{k}^{3}I$ and $\partial_{k}^{3}E_{W}$ to isolate this term after lower-order $k$-independent pieces cancel (Appendix D). The physical significance is clear: $P_{S}$ and $P_{IE}$ quantify how the UV geometry deforms in response to the tuning parameter $k$, encoding information about the RG flow towards different IR fixed points.
This UV sensitivity offers a unique advantage. While IR entanglement measures often mix quantum and classical (thermal) correlations, UV measures primarily probe quantum correlations because they capture short-range scales where quantum fluctuations dominate over thermal effects. This is evident from the AdS/CFT mapping: UV observables probe the near-boundary geometry ($z\to 0$), far from the thermal horizon at $z=1$, thus isolating quantum entanglement from thermal contributions that dominate at larger (IR) scales. This allows for a cleaner extraction of quantum entanglement signatures of QPTs, even at finite temperatures where these UV diagnostics remain robust. The universality of these UV signatures, confirmed across different model parameters ($\gamma$, $\lambda$) within the EMDA class and verified in the Q-lattice model [19], suggests they are a general feature of holographic MITs, rooted in the UV structure’s response to critical phenomena. This universality appears to stem from the generic way UV geometry deforms near critical points in holographic models, rather than depending on specific IR physics.
Robustness of UV Signatures at Finite Temperatures.— A crucial aspect of the UV diagnostics is their robustness against thermal fluctuations, particularly when contrasted with IR. We find that the sharp extrema observed in the derivatives of HEE, MI, EWCS, and conductivity in the UV region persist with essentially unchanged magnitude and location across a range of low, but not necessarily ultra-low, temperatures. More importantly, across a range of temperatures $T=0.002$ to $0.015$, the UV observables (e.g., the extremum in $\partial_{k}S$) exhibit smaller deviations from the zero-temperature critical point $k_{c_{0}}$ than the DC conductivity, and converge toward the critical point more rapidly than their DC conductivity as the temperature decreases, as shown in Fig. 4.
Figure 4: Absolute deviation of the critical value $k_{c}$, determined by various phase diagnostic quantities, from the zero-temperature limit reference $k_{c_{0}}$, as a function of $T$.
We can understand this temperature insensitivity in the UV region from the AdS/CFT correspondence. UV observables probe the asymptotic region of the bulk geometry ($z\to 0$), which lies far from the IR region and is thus inherently less affected by the presence and properties of the black brane horizon (located at $z=1$, corresponding to finite temperature $T>0$). From another perspective, Quantum fluctuations primarily govern short-distance correlations, which dominate in this near-boundary region, rather than thermal effects that become more significant deeper in the bulk and at larger subsystem sizes (IR).
Conversely, thermal entropy usually dominates IR observables at finite temperature. While thermal entropy itself can exhibit changes across a phase transition, it primarily reflects classical statistical correlations and can obscure the purely quantum entanglement features of a QPT. The UV region, being relatively decoupled from these dominant thermal contributions, provides a clearer window into the quantum critical aspects. This UV characteristic has significant practical implication: the UV diagnostics identified in this work reveal signatures of quantum criticality at temperatures considerably higher than those typically assumed necessary for IR-based probes. This is particularly advantageous for systems where achieving the extremely low temperatures traditionally required to isolate quantum critical behavior is experimentally challenging. This opens new avenues for investigating quantum critical phenomena, potentially simplifying experimental requirements and broadening the range of systems where we can effectively diagnose QPTs.
Summary and discussion.— In this letter, we have demonstrated that UV observables in holographic systems can serve as potent diagnostics for QPTs, challenging the conventional wisdom that such phenomena are exclusively governed by IR physics. Table 1 presents a summary of the distinct physical roles and characteristics of UV and IR observables. Focusing on a holographic MIT, we have revealed that derivatives of UV-dominated quantities—specifically, the first derivative of HEE and the third derivatives of MI, EWCS, and high-frequency AC conductivity—exhibit extrema near the QCP.
These UV signatures are not coincidental. They arise from the leading $k$-dependent deviations from the asymptotic AdS geometry, as shown by our analytical expansions in Appendix D. This implies that the UV region, often considered non-universal, encodes crucial information about the underlying RG flow and the impending phase transition. A key advantage of these UV probes is their robustness at finite temperature, where they continue to provide clear diagnostics even when IR signatures might be obscured by thermal effects. This also allows for a cleaner distinction between quantum and classical correlations associated with the QPT.
The robustness of these UV diagnostics is notable. Our analytical UV expansions are largely model-independent, not relying on specific coupling parameters like $\gamma$. Numerically, the extremal UV behaviors near the QCP persist for different choices of $\gamma$ and $\lambda$. Crucially, these UV signatures of criticality are not confined to the EMDA model; we have verified their presence in other frameworks, such as the Q-lattice model [19], indicating they represent a general feature of the UV region in holographic QPTs.
Our findings establish a new paradigm for investigating QPTs, shifting focus towards the rich, yet often overlooked, information content of the UV sector. This fundamentally reshapes our understanding of the interplay between UV and IR dynamics in holographic systems and quantum criticality. The ability to detect QPTs from short-distance physics opens new avenues for theoretical exploration, potentially extending to topological QPTs or systems beyond the Landau-Ginzburg framework. Furthermore, it may inspire novel experimental probes sensitive to UV correlations in strongly correlated materials. We anticipate that this IR-UV dual diagnostic framework could be extended to a wider class of physical quantities and models, offering a more complete picture of critical phenomena.
Table 1: Comparison between IR and UV observables.
IR observable UV observable
Probe low-energy excitations Probe high-energy degrees of freedom
Capture long-range correlations Capture short-range correlations
Limit: Smearing of the thermal signal Advantage: Distinction of the critical signal
Acknowledgement.— Peng Liu would like to thank Yun-Ha Zha and Yi-Er Liu for their kind encouragement during this work. This work is supported by the Natural Science Foundation of China under Grant Nos. 12475054, 12035016, 12275275, 12375055 and the Guangdong Basic and Applied Basic Research Foundation No. 2025A1515012063. Zhe Yang is supported by the Jiangsu Postgraduate Research and Practice Innovation Program under Grant No. KYCX25_3922.
Appendix A Boundary conditions and Hawking temperature
To obtain the background solution numerically, we impose the following UV boundary conditions:
$U(0)=1,\,V_{1}(0)=1,\,V_{2}(0)=1,\,a(0)=1,\,\phi(0)=\bar{\lambda},$ (5)
where $\bar{\lambda}$ is the strength of the lattice deformation. In addition, Hawking temperature follows the relation
$\bar{T}=\frac{12-\mu^{2}}{16\pi}U(1).$ (6)
Appendix B Details of the AC Conductivity Calculation
To compute the AC conductivity along the direction of lattice, i.e., $x$-direction, we introduce linear perturbations of the form
$g_{tx}=e^{-i\omega t}\delta h_{tx}(z),\quad A_{x}=e^{-i\omega t}\delta a_{x}(z),\quad\chi=e^{-i\omega t}\delta\chi(z),$ (7)
Figure 5: The schematic of the solution $da_{x}$ of the high-frequency AC conductivity at $\omega/\mu=20$ and $T=0.005$. Upper panel: The solution of $da_{x}$ and it shows rapid oscillations. Lower panel: The residual of the solution $da_{x}$, which demonstrate our solution is highly reliable.
leading to a set of coupled linear differential equations for $\{\delta h_{tx}(z),\delta a_{x}(z),\delta\chi(z)\}$. At the horizon $z=1$, we impose regular and ingoing boundary conditions
$\delta h_{tx}(z)$ $=(1-z)^{1-\frac{i\omega}{4\pi\mu T}}dh_{tx}(z),$ (8)
$\delta a_{x}(z)$ $=(1-z)^{-\frac{i\omega}{4\pi\mu T}}da_{x}(z),$
$\delta\chi(z)$ $=(1-z)^{-\frac{i\omega}{4\pi\mu T}}d\chi(z),$
where $dh_{tx}(z)$, $da_{x}(z)$, and $d\chi(z)$ are regular at the horizon.
At the UV boundary $z=0$, we set $da_{x}(0)=1$ to define the source of the external electric field. Additionally, to ensure that the response corresponds to a current-current correlator in the dual field theory, we impose the constraint
$d\chi(0)-\frac{ik}{\omega}dh_{tx}(0)=0,$ (9)
which comes from the diffeomorphism and gauge invariance.
Solving these equations numerically yields the perturbations $\{\delta a_{x}(z),\delta h_{tx}(z),\delta\chi(z)\}$, and finally, the AC conductivity is extracted via the Kubo formula:
$\sigma(\omega)=\frac{\delta a_{x}^{\prime}(z)}{i\omega\delta a_{x}(z)}\Bigg{|}_{z\to 0}.$ (10)
Our results show robust numerical convergence. All numerical calculations have been carefully validated and double-checked to ensure the reliability of our findings.
Two main numerical challenges arise when we evaluate $\sigma_{AC}$ at low temperatures and high frequencies:
  1. 1.
    Rapid oscillations: The perturbations $\{da_{x}(z),dh_{tx}(z),d\chi(z)\}$ exhibit strong oscillatory behavior near the horizon. This requires a high spatial resolution along the radial direction $z$ to accurately capture the rapid oscillatory behavior of perturbations, as shown in Fig. 5. For example, in the case $T=0.005$ and $\omega/\mu=20$, we discretize the domain using $1000$ grid points in the $z-$direction to ensure numerical stability and convergence. Since the oscillation scale is controlled by the ratio $\omega/T$, increasing $\omega/T$ demands higher spatial resolution.
  2. 2.
    Small signal strength: In the UV limit, the real part of the AC conductivity approaches $1$ asymptotically. The critical signal arises from subtle $k$-dependent deviations from this limiting value. As shown in Table 2, for fixed values of $(T,\,\omega/\mu)$, the variation in $\mathrm{Re}[\sigma_{AC}]$ across the interval $k\in[0.38,0.53]$ is extremely small—on the order of $10^{-6}$ to $10^{-8}$—requiring high-precision numerics to identify critical signal. This subtle signal stems from its appearance at the order $\omega^{-4}$ in the high-frequency expansion, as shown in Eq. 3.
$k$
$(T,\ \omega/\mu)$ 0.33 0.38 0.43 0.48 0.53
(0.01, 20) 1.00126150 1.00125966 1.00125760 1.00125533 1.00125267
(0.002, 20) 1.00126148 1.00125963 1.00125756 1.00125528 1.00125262
(0.01, 50) 1.00020052 1.00020050 1.00020047 1.00020045 1.00020041
Table 2: Real part of the high-frequency conductivity $\mathrm{Re}[\sigma_{AC}]$ for varying $k$ at different $(T,\omega/\mu)$. For clarity, only five representative $k$-values are shown here. In our calculations, the full interval $k\in[0.33,0.53]$ is discretized with 54 points.
Appendix C Holography quantum information
EE is a key measure of quantum entanglement, particularly suitable for pure states. The area of the minimal surface in the bulk gives its holographic counterpart, HEE [29, 30, 31, 32]:
$S_{A}=\frac{\text{Area}(\gamma_{A})}{4G_{N}},$ (11)
where $S_{A}$ is the EE, $G_{N}$ is Newton’s constant, and $\gamma_{A}$ is the minimal surface.
We examine a strip that is infinitely long in the $y$-direction and has a finite width $w$ in the $x$-direction, as depicted in Fig. 6. This allows us to focus on the $x$-direction where $\chi$ breaks translational symmetry. The HEE is characterized by the maximum depth $z_{s}$ of the minimal surface. For the metric in Eq. (2), $w$ and the HEE $S$ are related to $z_{s}$ as:
[…]
$S$ $=\frac{L_{y}}{2\mu G_{N}}\left[-\frac{1}{z_{s}}+\int_{0}^{z_{s}}\frac{dz}{z^{2}}\left(\frac{z_{s}^{2}V_{1}(z)V_{2}(z)}{\sqrt{P(z)V_{1}(z)W(z_{s},z)}}-1\right)\right],$
where $W(z_{s},z)\equiv z_{s}^{4}V_{1}(z)V_{2}(z)-z^{4}V_{1}(z_{s})V_{2}(z_{s})$, and $L_{y}(\to\infty)$ is the length of the infinitely long strip in the $y$ direction.
Quantum states in real systems are typically mixed states, for which HEE is insufficient as an entanglement measure. To quantify entanglement in mixed states, MI has been proposed. For disjoint regions $A$ and $C$ separated by $B$, MI is defined as [33, 34]:
$I(A;\,C)=S(A)+S(C)-S(A\cup C),$ (13)
where $S(X)$ is the HEE of region $X$. The HEE of the combined region $A\cup C$ is given by:
$S(A\cup C)=\min\{S(A)+S(C),\,S(B)+S(A+B+C)\}.$ (14)
MI provides a clear indicator of entanglement: $I(A;\,C)=0$ for non-entangled regions, and $I(A;\,C)>0$ for entangled regions.
Figure 6: Left: Schematic plot of HEE, where the red surface represents the geometry of HEE. Right: Schematic plot of EWCS in $x-z$ plane. The red curve shows the minimum cross-section between minimum surfaces $C_{1}$ and $C_{2}$ of the entanglement wedge. The blue and purple arrows represent tangent vectors $\left(\frac{\partial}{\partial z}\right)^{a}$ and $\left(\frac{\partial}{\partial\theta}\right)^{a}$, respectively. The local perpendicular condition of these tangent vectors determines the position of EWCS. The dark horizontal line represents the black brane horizon.
Another proposed measure for mixed state entanglement is the EWCS. The EWCS is defined as the area of the minimum cross-section within the entanglement wedge [35]: $E_{W}(\rho_{AB})=\min_{\Sigma_{AB}}\left(\frac{\text{Area}(\Sigma_{AB})}{4G_{N}}\right).$ The EWCS is calculated as:
[…]
with the corresponding equation of motion:
$x^{\prime}(z)^{3}\left(\frac{g_{xx}g^{\prime}_{yy}}{2g_{yy}g_{zz}}+\frac{g^{\prime}_{xx}}{2g_{zz}}\right)+x^{\prime}(z)\left(\frac{g^{\prime}_{xx}}{g_{xx}}+\frac{g^{\prime}_{yy}}{2g_{yy}}-\frac{g^{\prime}_{zz}}{2g_{zz}}\right)+x^{\prime\prime}(z)=0,$ (16)
where primes denote derivatives with respect to $z$, and $C_{p_{1},p_{2}}$ is the global minimal curve connecting the minimum surfaces $C_{1}$ and $C_{2}$, as depicted in Fig. 6. For HEE calculations, the configuration parameter $w$ represents the width of the single region under consideration. For MI and EWCS calculations, the configuration parameters $(a,b,c)$ represent the widths of regions $A$, $B$, and $C$, respectively.
In addition to UV observables, we also examine the MI and EWCS in the IR region, both of which exhibit extrema near the QCP, as shown in Fig. 7.
Figure 7: The schematic of MI (a) and first derivative of EWCS (b) with respect to $k$ for $a=c=40$, with $b$ varied. The black dots indicate the extrema of each curve, and the black dashed line represents the critical point.
Appendix D UV Series Expansion of Quantum Entanglement Measures
Here we detail the UV series expansion for HEE, MI, and EWCS. For HEE, we consider a strip of width $w$ and length $L_{y}$ (taken to infinity). The HEE $S$ and width $w$ are determined by the turning point $z_{s}$ of the minimal surface in the bulk, as given by Eq. (12).
In the UV limit, $w\to 0$, which corresponds to $z_{s}\to 0$. Expanding $w(z_{s})$ and $S(z_{s})$ from Eq. (12) in powers of $z_{s}$:
$w$ $=2\mu\left[\Gamma z_{s}+f_{1}z_{s}^{2}+f_{2}z_{s}^{3}+f_{3}z_{s}^{4}+O(z_{s}^{5})\right],$ (17)
$S$ $=\frac{L_{y}}{2\mu G_{N}}\left[-\frac{\Gamma}{z_{s}}+g_{1}+g_{2}z_{s}+g_{3}z_{s}^{2}+O(z_{s}^{3})\right],$
where $\Gamma=\frac{\sqrt{\pi}\Gamma(3/4)}{\Gamma(1/4)}$ is a constant. The coefficients $f_{i}$ and $g_{i}$ are functions of the near-boundary values of the metric and scalar field components and their $z$-derivatives at $z=0$: $\hat{U}^{(n)}\equiv\partial_{z}^{n}U|_{z=0}$, $\hat{V}_{1}^{(n)}\equiv\partial_{z}^{n}V_{1}|_{z=0}$, etc. These are related by the equations of motion at the boundary, e.g.,
$\hat{U}^{(1)}=\hat{V}_{1}^{(1)}=\hat{V}_{2}^{(1)},\quad\hat{U}^{(2)}=\hat{V}_{1}^{(2)}=\hat{V}_{2}^{(2)},$ (18)
$\left(\hat{U}^{(1)}\right)^{2}-2\hat{U}^{(2)}=3\left(\hat{\phi}^{(0)}\right)^{2},$
$\hat{V}_{2}^{(3)}+\hat{V}_{1}^{(3)}=-6\left(\hat{\phi}^{(0)}\right)^{2}\hat{V}_{1}^{(1)}-12\hat{\phi}^{(0)}\hat{\phi}^{(1)},$
where $\hat{\phi}^{(0)}=\mu\lambda$. The explicit forms of Eq. (17) are:
$\lim_{z_{s}\to 0}w=2\mu\Bigg{[}\Gamma z_{s}-\frac{1}{2}\hat{V}_{1}^{(1)}\Gamma z_{s}^{2}+\left(\frac{\Gamma}{4}\left(\hat{V}_{1}^{(1)}\right)^{2}-\frac{\pi-12\Gamma^{2}}{32\Gamma}\mu^{2}\lambda^{2}\right)z_{s}^{3}$ (19)
$+\left(\frac{(4+\mu^{2})\pi}{64}+\frac{3\pi-8\pi\Gamma-20\Gamma^{2}}{64\Gamma}\mu^{2}\lambda^{2}\hat{V}_{1}^{(1)}-\frac{\Gamma}{8}\left(\hat{V}_{1}^{(1)}\right)^{3}-\frac{\pi-2\Gamma}{4}\mu\lambda\hat{\phi}^{(1)}-\frac{\pi}{96}(\hat{U}^{(3)}+\hat{V}_{1}^{(3)})\right)z_{s}^{4}+O(z_{s}^{5})\Bigg{]},$
$\lim_{z_{s}\to 0}S=\frac{L_{y}}{2\mu G_{N}}\Bigg{[}-\frac{\Gamma}{z_{s}}+\frac{1-\Gamma}{2}\hat{V}_{1}^{(1)}+\frac{3(4\Gamma^{2}-\pi)}{32\Gamma}\mu^{2}\lambda^{2}z_{s}$
$+\Bigg{(}\frac{\pi(4+\mu^{2})}{32}+\frac{3\pi-16\pi\Gamma+4\Gamma^{2}}{64\Gamma}\mu^{2}\lambda^{2}\hat{V}_{1}^{(1)}+\frac{\Gamma-\pi}{2}\mu\lambda\hat{\phi}^{(1)}-\frac{\pi}{48}\left(\hat{U}^{(3)}+\hat{V}_{1}^{(3)}\right)\Bigg{)}z_{s}^{2}+O(z_{s}^{3})\Bigg{]}.$
Inverting $w(z_{s})$ to find $z_{s}(w)$:
$z_{s}=\frac{w}{2\Gamma\mu}+\frac{\hat{V}_{1}^{(1)}w^{2}}{8\Gamma^{2}\mu^{2}}+\left[(\pi-12\Gamma^{2})\mu^{2}\lambda^{2}+8\Gamma^{2}\left(\hat{V}_{1}^{(1)}\right)^{2}\right]\frac{w^{3}}{256\Gamma^{5}\mu^{3}}+\bigg{[}-3\Gamma\pi(4+\mu^{2})$ (20)
$+6\left(\pi+4\pi\Gamma-20\Gamma^{2}\right)\mu^{2}\lambda^{2}\hat{V}_{1}^{(1)}+24\Gamma^{2}\left(\hat{V}_{1}^{(1)}\right)^{3}+48\Gamma(\pi-2\Gamma)\mu\lambda\hat{\phi}^{(1)}+2\pi\Gamma\left(\hat{U}^{(3)}+\hat{V}_{1}^{(3)}\right)\bigg{]}\frac{w^{4}}{3072\Gamma^{6}\mu^{4}}+O(w^{5}).$
Substituting $z_{s}(w)$ into $S(z_{s})$ yields $S(w)$ for small $w$:
$S(w)=\frac{L_{y}}{4\mu G_{N}}\bigg{[}-\frac{4\Gamma^{2}\mu}{w}+\underbrace{\hat{V}_{1}^{(1)}}_{P_{S}}-\frac{\pi\lambda^{2}\mu}{16\Gamma^{2}}w+\left(C_{21}+C_{22}\underbrace{(\hat{U}^{(3)}-\hat{V}_{1}^{(3)}-2\hat{V}_{2}^{(3)})}_{P_{IE}}\right)w^{2}+O(w^{3})\bigg{]},$ (21)
where $C_{21}=\frac{\pi(4+\mu^{2})}{128\Gamma^{2}\mu^{2}}$ and $C_{22}=-\frac{\pi}{192\Gamma^{2}\mu^{2}}$ are constants independent of $k$. The $k$-dependence of $S(w)$ at leading order (constant term in $w$) comes from $P_{S}\equiv\hat{V}_{1}^{(1)}$. The next significant $k$-dependent term appears at $O(w^{2})$ through $P_{IE}\equiv\hat{U}^{(3)}-\hat{V}_{1}^{(3)}-2\hat{V}_{2}^{(3)}$.
For MI and EWCS, we consider a symmetric configuration $a=c=n\,b$ (for some constant $n$) and take $b\to 0$. In the presence of entanglement, combining Eqs. (13), (14) and the UV expansion of $S$ (21) yields:
$I=S(nb)+S(nb)-S((2n+1)b)-S(b)=\frac{L_{y}}{4\mu G_{N}}\bigg{[}C_{I0}\frac{1}{b}+C_{I1}b+\left(C_{I21}+C_{I22}P_{IE}\right)b^{2}+O(b^{3})\bigg{]},$ (22)
where $C_{I0},C_{I1},C_{I21}$, and $C_{I22}$ are $k$-independent coefficients derived from the $w^{-1},w^{1}$, and $w^{2}$ terms in Eq. (21). Importantly, the $P_{S}$ term cancels out in MI, and the leading $k$-dependence of MI comes from $P_{IE}$ at $O(b^{2})$.
For EWCS, in the symmetric configuration $a=c$, the minimal cross-section $\Sigma_{AB}$ is a surface at constant $x$ that extends from $z=0$ to some $z_{s}=z_{max}$). In the UV limit ($b\to 0$), the dominant contribution comes from $z\approx 0$. A similar expansion for $E_{W}$ (Eq. (15)) yields:
$E_{W}=\frac{L_{y}}{4\mu G_{N}}\bigg{[}C_{E0}\frac{1}{b}+C_{E1}b+\left(C_{E21}+C_{E22}P_{IE}\right)b^{2}+O(b^{3})\bigg{]}.$ (23)
Again, $C_{E0},C_{E1},C_{E21}$, and $C_{E22}$ are $k$-independent, and the leading $k$-dependence arises from $P_{IE}$ at $O(b^{2})$. The precise coefficients $C_{I0},C_{I1},C_{I21},C_{I22}$ and $C_{E0},C_{E1},C_{E21},C_{E22}$ depend on $n$ and other constants but their $k$-independence is what matters.
Thus, in the UV limit, both MI and EWCS have their leading $k$-dependence governed by the same geometric deformation term $P_{IE}\equiv\hat{U}^{(3)}-\hat{V}_{1}^{(3)}-2\hat{V}_{2}^{(3)}$:
$\partial_{k}^{n}I\propto\partial_{k}^{n}E_{W}\propto\partial_{k}^{n}P_{IE}.$ (24)
This explains why their third derivatives with respect to $k$ exhibit similar extremal behavior near the QCP, as shown in Fig. 3.